OpenStudy (anonymous):

Who can answer this? Medals given:

OpenStudy (anonymous):

If 3r = 4s, then r/s = ?

OpenStudy (blurbendy):

3r / s = 4

OpenStudy (anonymous):

4/3

OpenStudy (blurbendy):

3r / s = 4 1/3 *3r/s = 4*1/3 r/s = 4/3